Can you help me please..........thankyou
Terry is selling plants from her garden. Buy 3/$4.50 and get one free.
i)Mrs. Greenthumb wants 20 plants; how much will she pay?
ii)How many plants would you get for $27???

Answer by vertciel(183)

Hello there,
I divided your question into two parts.
i) Since Terry is selling 3 flowers + 1 free flower for $4.5, the price can actually be denoted as 4 flowers/$4.50. Do you see why? The flower is free so I add 1 to the price without adding more money to $4.5.
Now, Mrs. Greenthumb wants 20 flowers so you write up a ratio:
4/4.50 = 20/x <-- x represents the amount Mrs. Greenthumb pays.
Find x and you have your answer.

ii) Similar to part i), you write another ratio based on the fact that you get 4 flowers for $4.5.
4/4.5 = x/$27 <-- x represents the number of flowers Mrs. Greenthumb receives. Note that x is now on the top (numerator).

Find x and you have your answer.